OpenStudy (anonymous):

Use the quadratic formula for (t^2 - 3)^2.

OpenStudy (anonymous):

I got the answer in decimals by factoring by my teacher said it has to be in radical form.

OpenStudy (zarkon):

are you trying to find the roots of \[(t^2-3)^2\] i.e. finding the solutions to \[(t^2-3)^2=0\] ?

OpenStudy (anonymous):

Yes, the zeros (roots).

OpenStudy (zarkon):

ok \[(t^2-3)^2=0\] take square root \[(t^2-3)=\sqrt{0}=0\] add 3 \[t^2=3\] take square root again \[t=\pm\sqrt{3}\]

OpenStudy (anonymous):

And what would be the two zeros (or roots)? Can you go any farther and not have decimals?

OpenStudy (zarkon):

that is it \[t=\sqrt{3}\] and \[t=-\sqrt{3}\]

OpenStudy (anonymous):

Oh! Wow, I can't believe I never understood that. Thank you, you just explained a bunch to me.

OpenStudy (zarkon):

no problem